# REBEL: Reinforcement Learning via Regressing Relative Rewards

### v1

@@ 7 Summary and Future Work
In summary, we propose REBEL, an RL algorithm that reduces the problem of RL to solving a sequence of relative reward regression problems on iteratively collected datasets. In contrast to policy gradient approaches that require additional networks and heuristics like clipping to ensure optimization stability, REBEL requires that we can drive down training error on a least squares problem. This makes it strikingly simple to implement and scale. In theory, REBEL matches the best guarantees we have for RL algorithms in the agnostic setting, while in practice, REBEL is able to match and sometimes outperform methods that are far more complex to implement or expensive to run across both language modeling and guided image generation tasks.
@@
There are several open questions raised by our work. The first is whether using a loss function other than square loss (e.g. log loss or cross-entropy) could lead to better performance in practice (Farebrother et al., 2024) or tighter bounds (e.g. first-order / gap-dependent) in theory (Foster and Krishnamurthy, 2021; Wang et al., 2023a, 2024). The second is whether, in the general (i.e. non-utility-based) preference setting, the coverage condition assumed in our analysis is necessary – we conjecture it is. Relatedly, it would be interesting to explore whether using preference (rather than reward) models to provide supervision for REBEL replicates the performance improvements reported by Swamy et al. (2024); Munos et al. (2023). Third, while we focus primarily on the bandit setting in the preceding sections, it would be interesting to consider the more general RL setting and explore how offline datasets can be used to improve the efficiency of policy optimization via techniques like resets (Bagnell et al., 2003; Ross and Bagnell, 2014; Swamy et al., 2023; Chang et al., 2023, 2024).
@@ Appendix A Proof of Claim 1
where the last equality uses the fact that cross terms from completing the square are zero. Now recall Eq. 13 which is a ordinarly least square regression problem. The minimum norm solution of the least square regression problem is:
@@
$\quad\quad-\mathbb{E}_{x,y\sim\pi_{t}y^{\prime}\sim\mu}\nabla_{\theta}\ln\pi_{\theta_{t}}(y^{\prime}|x)r(x,y)\Big{)}$
@@
Appendix B Proof of Theorem 1
@@
Appendix C Proof of Theorem 2
@@
Appendix D Additional Experiment Details
D.1 Dataset Details
We present dataset details in Table 3.
Table 3: Dataset split, prompts, and maximum generation length
@@ D.1 Dataset Details
D.2 Hyperparameter Details
@@
D.3 Winrate Details
Below we show the prompt for winrate evaluation and an example evaluation from GPT4.
@@ D.3 Winrate Details
Appendix E Example Generations
E.1 Summarization Generations
@@
E.2 Image Generations
Appendix F Ablation Analysis
@@
Table 4: REBEL ablation of the key hyperparameter $\eta$ The best-performing $\eta$ for each metric is highlighted in bold.
Just like DPO, tuning REBEL is much more straightforward than PPO since the only hyperparameter REBEL introduced is $\eta$. We investigate how sensitive REBEL is to learning rate $\eta$ in the loss. The results of ablation is shown in Table 4 with the same setting detailed in Appendix D.2 except for $\eta$. REBEL achieves the best performance when $\eta=1$, while increasing or decreasing $\eta$ leads to decreased performance. Our result here indicates that $\eta$ is an important hyperparameter that requires tuning for achieving a good performance.
Appendix G Regression Loss During Training
Figure 6: REBEL’s reward difference prediction error throughout training of our 6.9B parameter policy on the TL;DR task. The reward used for this task is unbounded with the range of values of the human labels in the validation set being $[-6.81,7.31]$. We plot both the smoothed values with a moving average and the loss vales at each iteration.
Figure 6 shows the observed loss or Eq. 9 that we observed when finetuning the 6.9B Pythia model on TL;DR. We see that REBEL minimizes the loss throughout training maintaining a relatively low mean squared error given that our observed rewards were mostly between $[-10,10]$. Note that our learned reward model, however, is unbounded.

### v2

@@ 7 Conclusion and Future Work
We propose REBEL, an RL algorithm that reduces the problem of RL to solving a sequence of relative reward regression problems on iteratively collected datasets. In contrast to policy gradient approaches that require additional networks and heuristics like clipping to ensure optimization stability, it suffices for REBEL to merely drive down error on a least squares problem, making it strikingly simple to implement and scale. In theory, REBEL matches the best guarantees we have for RL algorithms in the agnostic setting, while in practice, REBEL is able to match and sometimes outperform methods that are far more complex to implement or expensive to run across both language modeling and guided image generation tasks.
@@
There are several open questions raised by our work. The first is whether using a loss function other than square loss (e.g. log loss or cross-entropy) could lead to better performance in practice (Farebrother et al., 2024) or tighter bounds (e.g. first-order / gap-dependent) in theory (Foster and Krishnamurthy, 2021; Wang et al., 2023, 2024). The second is whether, in the general (i.e. non-utility-based) preference setting, the coverage condition assumed in our analysis is necessary – we conjecture it is. Relatedly, it would be interesting to explore whether using preference (rather than reward) models to provide supervision for REBEL replicates the performance improvements reported by Swamy et al. (2024); Munos et al. (2023). Third, while we focus primarily on the bandit setting in the preceding sections, it would be interesting to consider the more general RL setting and explore how offline datasets can be used to improve the efficiency of policy optimization via techniques like resets (Bagnell et al., 2003; Ross and Bagnell, 2014; Swamy et al., 2023; Chang et al., 2023, 2024).
Acknowledgements
WS acknowledges funding from NSF IIS-2154711 and NSF CAREER 2339395; JDL acknowledges support of the NSF CCF 2002272, NSF IIS 2107304, and NSF CAREER Award 2144994. Jonathan Chang is supported by LinkedIn under the LinkedIn-Cornell Grant. Kiante Brantley is supported by NSF under grant No. 2127309 to the Computing Research Association for the CIFellows Project.
@@ Appendix A Proof of Claim 1
where the last equality uses the fact that cross terms from completing the square are zero. Now recall Eq. 9 which is an ordinarly least square regression problem. The minimum norm solution of the least square regression problem is:
@@
$\quad\quad-\mathbb{E}_{x,y\sim\pi_{t},y^{\prime}\sim\mu}\nabla_{\theta}\ln\pi_{\theta_{t}}(y^{\prime}|x)r(x,y)\Big{)}$
@@
Appendix B Proof of Claim 2
We prove claim 2 in this section. We start by approximating our predictor $\frac{1}{\eta}\ln\pi_{\theta}(y|x)/\pi_{\theta_{t}}(y|x)$ by its first order Taylor expansion at $\theta_{t}$: $\frac{1}{\eta}\left(\ln\pi_{\theta}(y|x)-\ln\pi_{\theta_{t}}(y|x)\right)\approx\frac{1}{\eta}\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)^{\top}(\theta-\theta_{t})$, where $\approx$ indicates that we ignore higher order terms in the expansion. Setting $\delta:=\theta-\theta_{t}$ and replace $\frac{1}{\eta}\left(\ln\pi_{\theta}(y|x)-\ln\pi_{\theta_{t}}(y|x)\right)$ by its first order approximation in Eq. 1, we arrive at:
$\min_{\delta}\sum_{(x,y,y^{\prime})\in\mathcal{D}_{t}}\left(\frac{1}{\eta}\left(\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)-\nabla_{\theta}\ln\pi_{\theta_{t}}(y^{\prime}|x)\right)^{\top}\delta-\left(r(x,y)-r(x,y^{\prime})\right)\right)^{2}$ (11)
under finite setting.
Following the previous derivation, we have the unbiased estimate of Fisher information matrix under the finite setting as:
$\hat{F}_{t}$ $:=\frac{2}{\eta^{2}N}\sum_{x_{n},y_{n}\sim\pi_{mix}}\nabla_{\theta}\ln\pi_{\theta_{t}}(y_{n}|x_{n})\nabla_{\theta}\ln\pi_{\theta_{t}}(y_{n}|x_{n})^{\top}$
Since Eq. 11 is an ordinarly least square regression problem. The minimum norm solution of the least square regression problem is:
$\delta$ $=(\eta/2)\tilde{\hat{F}}_{t}^{\dagger}\frac{1}{N}\sum_{n}\left(\nabla_{\theta}\ln\pi_{\theta_{t}}(y_{n}|x_{n})-\nabla_{\theta}\ln\pi_{\theta_{t}}(y^{\prime}_{n}|x_{n})\right)\left(r(x_{n},y_{n})-r(x_{n},y^{\prime}_{n})\right)$
$=\eta\tilde{\hat{F}}_{t}^{\dagger}\frac{1}{2N}\sum_{n}\left(\nabla\ln\pi_{\theta_{t}}(y_{n}|x_{n})(r(x_{n},y_{n})-r(x_{n},y^{\prime}_{n}))+\nabla\ln\pi_{\theta_{t}}(y^{\prime}_{n}|x_{n})(r(x_{n},y^{\prime}_{n})-r(x_{n},y_{n}))\right)$
Appendix C Extending REBEL to General Preferences
C.1 A Game-Theoretic Perspective on Learning from Preferences
This implies that $(\pi_{\mathsf{MW}},\pi_{\mathsf{MW}})$ is a Nash Equilibrium (Wang et al., 2023; Munos et al., 2023; Swamy et al., 2024; Ye et al., 2024). As is standard in game solving, our objective is to obtain an $\epsilon$-approximate MW $\widehat{\pi}$ measured by the duality gap (DG):
C.2 Self-Play Preference Optimization (SPO) with REBEL as Base Learner
Appendix D Proof of Theorem 1
@@
Appendix E Extension of analysis to General Preferences
Extending the above analysis to the general preference case is straightforward except that it requires a stronger coverage condition. This is because we want to find a Nash Equilibrium, which requires a comparison between the learned policy against all the other policies. Results from the Markov Game literature (Cui and Du, 2022; Zhong et al., 2022; Cui and Du, 2022; Xiong et al., 2023) and Cui and Du (2022) have shown that the standard single policy coverage condition used in single-player optimization is provably not sufficient. In particular, they propose using a notion of unilateral concentrability for efficient learning, which can be defined as
Over $T$ iterations, assume that for all $t$, we have:
for some $\epsilon$.
Theorem 2.
Note that the coverage condition here is much stronger than the single policy coverage condition in the RL setting. We conjecture that this is the cost one has to pay by moving to the more general preference setting and leaving the investigation of the necessarily coverage condition for future work.
E.1 Proof of Theorem 2
@@
Appendix F Additional Experiment Details
F.1 Summarization
F.1.1 Dataset Details
We present dataset details in Table 4. Dataset available at https://github.com/openai/summarize-from-feedback
Table 4: Dataset split, prompts, and maximum generation length for TL;DR summarization
@@ F.1.1 Dataset Details
F.1.2 Model Details
For SFT models, we train a Pythia 1.4B (Biderman et al., 2023) [^1] HuggingFace Model Card: EleutherAI/pythia-1.4b-deduped model for 1 epoch over the dataset with human references as labels, and use the existing fine-tuned 2.8B [^2] HuggingFace Model Card: vwxyzjn/EleutherAI_pythia-2.8b-deduped__sft__tldr and 6.9B [^3] HuggingFace Model Card: vwxyzjn/EleutherAI_pythia-6.9b-deduped__sft__tldr models. For reward models, we train a Pythia 1.4B parameter model for 1 epoch over the preference dataset and use the existing reward models with 2.8B [^4] HuggingFace Model Card: vwxyzjn/EleutherAI_pythia-2.8b-deduped__reward__tldr and 6.9B [^5] HuggingFace Model Card: vwxyzjn/EleutherAI_pythia-6.9b-deduped__reward__tldr parameters. For both REBEL and baseline methods using 1.4B and 2.8B parameters, we trained the policy and/or the critic using low-rank adapters (LoRA) (Hu et al., 2022) on top of our SFT and/or reward model respectively. For the 6.9B models, we perform full-parameter training. The 1.4B and 2.8B models are trained on 8 A6000 GPUs for one day and two days respectively. The 6.9B model is train on 8 H100 GPUs for two days.
F.1.3 Baseline Implementation Details
For supervised fine-tuning (SFT), reward modeling training, PPO, and DPO, we follow the implementation at https://github.com/vwxyzjn/summarize_from_feedback_details. For iterative dpo, we implement as follows:
Algorithm 2 Iterative DPO
1:  Input: Reward $r$, policy class $\Pi=\{\pi_{\theta}\}$, parameter $\beta$
2:  Initialize policy $\pi_{\theta_{0}}$.
3:  for $t=0$ to $T-1$ do
4:     Collect dataset $\mathcal{D}_{t}=\{x,y,y^{\prime}\}$ where $x\sim\rho,y\sim\pi_{t}(\cdot|x),y^{\prime}\sim\pi_{t}(\cdot|x)$
5:     Solve square loss regression problem:
$\theta_{t+1}=\mathop{\mathrm{argmin}}_{\theta}\sum_{(x,y,y^{\prime})\in\mathcal{D}_{t}}-\left[\ln\sigma\left((\beta\ln\frac{\pi_{\theta}(y|x)}{\pi_{\theta_{t}}(y|x)}-\beta\ln\frac{\pi_{\theta}(y^{\prime}|x)}{\pi_{\theta_{t}}(y^{\prime}|x)})sgn\left(r(x,y)-r(x,y^{\prime})\right)\right)\right]$ (21)
6:  end for
where $sgn$ is a sign function. Our implementation of iterative DPO is similar to REBEL where, at each iteration, we update with respect to $\pi_{\theta_{t}}$. The major difference is that REBEL regresses toward the differences in rewards while iterative DPO only utilizes the pairwise preference signal from the rewards.
F.1.4 Reward Details
To ensure that $\pi_{\theta}$ remains close to $\pi_{\theta_{0}}$, we apply an additional KL penalty to the reward:
$r(x,y)=RM(x,y)-\gamma(\ln\pi_{\theta_{t}}(y|x)-\ln\pi_{\theta_{0}}(y|x))$ (22)
where $RM(x,y)$ is score from the reward model given prompt $x$ and response $y$. Furthermore, to ensure that the generations terminate within the maximum generation length, we penalize any generation that exceeds this length by setting $r(x,y)$ to a small fixed constant, $\Gamma$.
For TL;DR summarization, we set $\gamma=0.05$ and $\Gamma=-1$.
F.1.5 Hyperparameter Details
@@
RLOO batch size: 512
learning rate: 3e-6
schedule: linear decay
train epochs: 1
kl coefficient: 0.05
K: 2 or 4
REINFORCE batch size: 512
learning rate: 3e-6
schedule: linear decay
train epochs: 1
kl coefficient: 0.05
Iterative DPO batch size: 512
learning rate: 3e-6
schedule: linear decay
train epochs: 1
num epochs: 4
$\beta$: 0.05
kl coefficient: 0.05
@@
F.1.6 Winrate Details
We are using gpt-4-0613 checkpoint for winrate computations. Below we show the prompt for winrate evaluation and an example evaluation from GPT4.
@@ F.1.6 Winrate Details
F.2 General Chat
F.2.1 Dataset Details
We present dataset details in Table 5.
Table 5: Dataset details for General Chat
Dataset Size Prompt Length Generation Length
Nectar [^6] HuggingFace Dataset Card: berkeley-nest/Nectar 183k 1024 1024
UltraFeedback [^7] HuggingFace Dataset Card: openbmb/UltraFeedback 64k 1024 1024
F.2.2 Model Details
For both OpenChat-3.5888HuggingFace Model Card: openchat/openchat_3.5 and Meta-Llama-3-8B-Instruct [^9] HuggingFace Model Card: meta-llama/Meta-Llama-3-8B-Instruct, we only train the last four layers in the model while keep other layers frozen. For both Starling-RM-7B-alpha [^10] HuggingFace Model Card: berkeley-nest/Starling-RM-7B-alpha and FsfairX-LLaMA3-RM-v0.1111111HuggingFace Model Card: sfairXC/FsfairX-LLaMA3-RM-v0.1, we directly use the reward scores without any normalizations. We filter out prompts that are longer than $1,024$ tokens ($2.3\%$) to fit the input length. Both experiments are trained on 8 H100 GPUs for four days.
F.2.3 Reward Details
To ensure that $\pi_{\theta}$ remains close to $\pi_{\theta_{0}}$, we apply an additional KL penalty to the reward:
$r(x,y)=RM(x,y)-\gamma(\ln\pi_{\theta_{t}}(y|x)-\ln\pi_{\theta_{0}}(y|x))$ (23)
where $RM(x,y)$ is score from the reward model given prompt $x$ and response $y$. Furthermore, to ensure that the generations terminate within the maximum generation length, we penalize any generation that exceeds this length by setting $r(x,y)$ to a small fixed constant, $\Gamma$.
For the general chat experiments, we set $\gamma=0.05$ and $\Gamma=-4$.
F.2.4 Hyperparameter Details
Parameter setting for General Chat Setting Parameters Setup 1
Base model: OpenChat-3.5
Reward Model: Starling-RM-7B-alpha
Dataset: Nectar
batch size: 32
learning rate: 1e-7
schedule: linear decay
train epochs: 1
num epochs: 4
$\eta$: 1.0
kl coefficient: 0.05
Setup 2
Base model: Meta-Llama-3-8B-Instruct
Reward Model: FsfairX-LLaMA3-RM-v0.1
Dataset: UltraFeedback
batch size: 32
learning rate: 1e-7
schedule: linear decay
train epochs: 1
num epochs: 4
$\eta$: 1.0
kl coefficient: 0.05
F.3 Image Generation
F.3.1 Dataset Details
Generation prompts: cat, dog, horse, monkey, rabbit, zebra, spider, bird, sheep, deer, cow, goat, lion, tiger, bear, raccoon, fox, wolf, lizard, beetle, ant, butterfly, fish, shark, whale, dolphin, squirrel, mouse, rat, snake, turtle, frog, chicken, duck, goose, bee, pig, turkey, fly, llama, camel, bat, gorilla, hedgehog, kangaroo.
F.3.2 Model Details
We use the latent consistency model (Luo et al., 2023) distillation of the Dreamshaper v7 model  [^12] Huggingface model card: SimianLuo/LCM_Dreamshaper_v7 for our experiments. Experiments are conducted on 4 A6000 GPUs with each run requiring 10 hours.
F.3.3 Hyperparameter Details
Parameter setting for Consistency Models Setting Parameters PPO advantage clip maximum: 10
batches per epoch: 10
clip range: 1e-4
learning rate: 1e-4
gradient accumulation steps: 8
max gradient norm: 5
number of epochs: 100
horizon: 8
number of sample inner epochs: 2
sample batch size (per GPU): 8
rolling statistics buffer size: 32
rolling statistics min count: 16
train batch size (per GPU): 2
LoRA rank: 8
Lora $\alpha$: 8
REBEL advantage clip maximum: 10
batches per epoch: 10
learning rate: 3e-4
$\eta$: 1
gradient accumulation steps: 8
max gradient norm: 5
number of epochs: 100
horizon: 8
number of sample inner epochs: 1
sample batch size (per GPU): 8
rolling statistics buffer size: 32
rolling statistics min count: 16
train batch size (per GPU): 1
LoRA rank: 8
Lora $\alpha$: 8
Appendix G Example Generations
G.1 Summarization Generations from a 2.8B model
@@ Appendix G Example Generations
G.2 Chat Generations from REBEL-Llama-3
Prompt Hi, I’d like to play ice hockey. Can you explain how the game works?
Generation Hockey enthusiast! I’d be happy to explain the basics of ice hockey. Here’s a comprehensive overview to get you started:
Objective:
The objective of ice hockey is to score more goals than the opposing team by using your stick to hit a puck (a rubber disc) into the opponent’s goal.
Gameplay:
1. Game Structure: A standard ice hockey game consists of three periods, each lasting 20 minutes.
2. Teams: Each team consists of six players on the ice at a time: Forwards (3): Players who score goals and create scoring opportunities. Defensemen (2): Players who defend their goal and prevent the opposing team from scoring. Goalie (1): The player responsible for preventing the puck from entering the goal. 3. Gameplay Flow: The game starts with a face-off, where two players from opposing teams try to gain control of the puck after it’s dropped onto the ice by the referee. Players can move the puck using their sticks, skates, or by carrying it.
4. Scoring: A goal is scored when a player hits the puck into the opponent’s goal using their stick, skate, or body.
5. Penalties: If a player breaks the rules, they can be sent to the penalty box for a set amount of time (2-5 minutes). During this time, their team must play shorthanded (with one fewer player on the ice).
6. Power Play: If the opposing team scores during a penalty, the penalized team gets a power play, where they have a man advantage on the ice.
7. Icing: If a player shoots the puck from behind their own blue line and it crosses the opponent’s goal line without being touched, it’s considered icing, and the opposing team is awarded a face-off in their own end zone.
8. Offside: A player cannot enter the opponent’s zone before the puck, or they’ll be called offside.
Common Penalties: Hooking: Using the stick to hook an opponent’s leg or body. Holding: Grabbing onto another player’s jersey or equipment. Tripping: Using a stick, leg, or body to trip an opponent. Interference: Interfering with an opponent who doesn’t have possession of the puck. Equipment: Stick: A long, flat stick used to pass, shoot, and control the puck. Skates: Specialized ice skates designed for speed, agility, and stopping. Pads: Protective gear worn by players, including shoulder pads, elbow pads, gloves, shin guards, and hockey pants. Helmet: A protective helmet worn by all players to prevent head injuries. Goalie Gear: Specialized equipment worn by goalies, including a large pad covering their body, gloves, and a mask. Basic Strategies: Forechecking: Applying pressure to the opponent’s defensemen to gain possession of the puck. Backchecking: Pressuring the opponent’s forwards to prevent them from scoring. Breakouts: Quickly moving the puck from your own end zone to the neutral zone to create scoring opportunities. This is a solid foundation to get you started. As you play more, you’ll learn more about the game’s nuances and strategies. Have fun playing ice hockey!
Prompt Why do people give Reddit Gold to admins?
Generation On Reddit, Reddit Gold (also known as G̈old)̈ is a virtual currency that allows users to show appreciation for content creators, moderators, and other users who contribute positively to the community. When someone gives Reddit Gold to an admin (or moderator), it’s often a way to express gratitude for their efforts in maintaining the community’s quality, resolving issues, and keeping the discussion civil.
Here are some reasons why people might give Reddit Gold to admins:
1. Appreciation for moderation: Admins work tirelessly to ensure that discussions remain on-topic, respectful, and follow community guidelines. Giving Gold acknowledges their dedication and hard work in maintaining a positive environment.
2. Recognition of expertise: Admins often possess in-depth knowledge about specific topics or communities. Giving Gold shows appreciation for their expertise and willingness to share it with others.
3. Support for community building: Admins play a crucial role in fostering engagement, encouraging participation, and creating opportunities for users to connect with each other. Gold gifts demonstrate support for their efforts in building a strong community.
4. Rewarding good behavior: When admins demonstrate excellent behavior, such as resolving conflicts fairly or providing helpful guidance, users might give Gold as a way to recognize and reward their good behavior.
5. Boosting morale: Giving Gold can be a morale booster for admins, acknowledging their value to the community and motivating them to continue their excellent work.
It’s worth noting that Reddit Gold is not required or expected, and admins don’t receive any direct benefits from receiving Gold. However, the gesture of appreciation can go a long way in fostering a positive and supportive community.
G.3 Image Generations
Example Generations of REBEL
Appendix H Ablation Analysis
@@
Table 6: REBEL ablation of the key hyperparameter $\eta$ on summarization task and 2.8B model. The best-performing $\eta$ for each metric is highlighted in bold.
Just like DPO, tuning REBEL is much more straightforward than PPO since the only hyperparameter REBEL introduced is $\eta$. We investigate how sensitive REBEL is to learning rate $\eta$ in the loss. The results of ablation on summarization task and 2.8B model is shown in Table 6 with the same setting detailed in Appendix F.1.5 except for $\eta$. REBEL achieves the best performance when $\eta=1$, while increasing or decreasing $\eta$ leads to decreased performance. Our result here indicates that $\eta$ is an important hyperparameter that requires tuning for achieving a good performance. Setting $\eta$ to $1.0$ is a good starting point since, for all of our experiments from language modeling to image generation, $\eta=1$ achieves the best results.
Appendix I Trade-off between Reward Model Score and KL-divergence
Figure 5: Plot of Reward vs KL-Divergence for 2.8B REBEL and PPO for summarization. We evaluate the models across the entire test set every 100 steps for 2,000 steps. Left: each point represents the average reward score and KL-divergence for a specific time step; the eclipse represents the confidence interval with 2 standard deviations. Right: we divide the KL distribution at the 2,000-step into $10$ bins with equal size and average the corresponding RM scores in each bin.
The trade-off between the reward model score and KL-divergence is shown in Figure 5. We evaluate the 2.8B REBEL and PPO every 400 gradient updates during training for 8,000 updates on summarization. The sample complexity of each update is held constant across both algorithms for fair comparison. For the left plot, each point represents the average divergence and score over the entire test set, and the eclipse represents the confidence interval with 2 standard deviations. As observed previously, PPO exhibits lower divergence, whereas REBEL shows higher divergence but is capable of achieving larger RM scores. Notably, towards the end of the training (going to the right part of the left plot), REBEL and PPO have similar KL and RM scores. For the right plot in Figure 5, we analyze a single checkpoint for each algorithm at the end of training. For each algorithm, we group every generation from the test set by its KL distribution into $10$ equally sized bins and calculate the average of the corresponding RM score for each bin. We can see that REBEL achieves higher RM scores for generations with small divergence while requiring larger divergence for generations with the highest scores.
Appendix J Regression Loss During Training
Figure 6: REBEL’s reward difference prediction error throughout training of our 6.9B parameter policy on the summarization task. The reward used for this task is unbounded with the range of values of the human labels in the validation set being $[-6.81,7.31]$. We plot both the smoothed values with a moving average and the loss vales at each iteration.
Figure 6 shows the observed loss of Eq. 1 that we observed when finetuning the 6.9B Pythia model on summarization. We see that REBEL minimizes the loss throughout training maintaining a relatively low mean squared error given that our observed rewards were mostly between $[-10,10]$. Note that our learned reward model, however, is unbounded.
Appendix K Breakdown of MT-Bench
Figure 7: Breakdown of MT-Bench results over eight dimensions.
Figure 7 shows the breakdown of MT-Bench results. REBEL-OpenChat-3.5 outperforms both Starling-LM-7B-alpha and OpenChat-3.5 on six out of eight dimensions including writing, roleplay, math, extraction, STEM, and humanities. REBEL-Llama-3 outperforms Meta-Llama-3-8B-Instruct on roleplay, reasoning, extraction, STEM, and humanities.